x+3y=11
3-y+2z=1
-2x=-4

Answer by amazingrace333(11)

Start where you only have one variable, in this case, "x".
-2x = -4, solve: x = -4/-2, so x=2
therefore, x+3y=11 becomes 2+3y=11. Solve: 3y=11-2, 3y = 9, therefore y=9/3 or 3
3-y+2z=1 becomes 3-3+2z=1. Using order of operations, 3-3 becomes 0 leaving 2z=1
solve for z by dividing both sides by 2, therefore z=1/2
Check your answers:
-2*2 = -4
2+3(3)=11
and finally
3-3+2(1/2)=3-3+1=0+1=1
So your answers are:
x=2, y=3 and z=1/2